OpenStudy (jjfun):

Solve |5x - 1| < 1 {x|0 < x < 2/5} {x|x < 0 or x > 2/5} {x|-2/5 < x < 2/5}

OpenStudy (tootsi123):

If you use mathway.com and type in this problem it will give you the answer and show you the work if desired.

OpenStudy (tootsi123):

Your answer is b

alones (alones):

5x−1|<1 Solve Absolute Value. |5x−1|<1 We know5x−1<1and5x−1>−1 5x−1<1(Condition 1) 5x−1+1<1+1(Add 1 to both sides) 5x<2 5x5<25(Divide both sides by 5) x<\(\frac{ 2}{ 5 }\) 5x−1>−1(Condition 2) 5x−1+1>−1+1(Add 1 to both sides) 5x>0 \(\frac{ 5x}{ 5 }\) \(\frac{ 5x}{ 5 }\)>\(\frac{ 0}{ 5 }\)(Divide both sides by 5) x>0 Answer: x< \(\frac{ 2}{ 5 }\) and x>0

alones (alones):

*\(\frac{ 2}{ 5 }\)

OpenStudy (sshayer):

\[\left| x-a \right|<b\] -b<x-a<b